OpenStudy (anonymous):

A submarine cuts through the water at an elevation of 203 meters under the sea. The sub rises, and the elevation changes by 4.3 meters per minute for the first 15 minutes. The sub continues to rise at 5.2 meters per minute for the next 12 minutes. What is the elevation of the submarine after rising? Assume that sea level is at an elevation of 0 feet. -126.9 m -76.1 m 126.9 m 339.9 m

OpenStudy (anonymous):

\[\large \sf -203+(4.3 \times 15)+(5.2 \times 12)=?\]